OpenStudy (anonymous):

how do you solve x+1/4 = 1/6 + 2-x/3 (the slashes mean over)

OpenStudy (anonymous):

it's (2-x) / 3 btw

OpenStudy (anonymous):

multiply the "/4" to the right side and the "/3" the the left side so 3(x+1) = 4(1/6 -x +2) 3x +3 = 2/3 -4x +8 7x = 2/3 + 5 7x = 17/3 x = (17/3)(1/7) x = 17/21

OpenStudy (anonymous):

hi